OpenStudy (freemap):

What is the distance between M(–4, 2) and N(6, –3)? Round to the nearest tenth.

jhonyy9 (jhonyy9):

your opinion please ?

OpenStudy (freemap):

whats the formula again

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\) does this help ??

OpenStudy (freemap):

yes

hartnn (hartnn):

what is the distance that you got?

OpenStudy (michele_laino):

hint: we can write this: \[\begin{gathered} d = \sqrt {{{\left( {{x_2} - {x_1}} \right)}^2} + {{\left( {{y_2} - {y_1}} \right)}^2}} = \hfill \\ \hfill \\ = \sqrt {{{\left( { - 4 - 6} \right)}^2} + {{\left( {2 - \left( { - 3} \right)} \right)}^2}} = \hfill \\ \hfill \\ = \sqrt {{{\left( { - 10} \right)}^2} + {{\left( 5 \right)}^2}} = ...? \hfill \\ \end{gathered} \]

OpenStudy (freemap):

-100+25 8.660

OpenStudy (michele_laino):

we have \((-10)^2=100\)

OpenStudy (freemap):

oh, yes sorry 100+25 11.180

OpenStudy (michele_laino):

that's right! Next please round such value to the nearest tenth

OpenStudy (freemap):

11.19

OpenStudy (michele_laino):

I think it is \(11.2\)

OpenStudy (freemap):

oh, yes I see

OpenStudy (freemap):

Thank you

OpenStudy (michele_laino):